### v1

# Jacobian rings and the infinitesimal Torelli Theorem

Abstract
In this article we deal with jacobian rings and identify a mixed Hodge component of a nondegenerate hypersurface in the torus with a lattice geometric quotient vector space. We introduce a period map, study its differential and compute the kernel of the differential much explicitly via certain Laurent polynomials. As an application we deal with the infinitesimal Torelli theorem.
1 Introduction
Being a subsequent paper to the paper ([Gie22]) summarizing, generalizing and extending results from my doctoral thesis ([Gie23]) we adopt the notation respectively. Given a nondegenerate Laurent polynomial
$f=\sum\limits_{m\in\Delta\cap M}a_{m}\cdot x^{m}$ (1)
with $n$-dimensional Newton polytope $\Delta$ take
$Z_{f}:=\{f=0\}\subset T.$
We are interested in the mixed Hodge components
$H^{n-k,k-1}H^{n-1}(Z_{f},\mathbb{C})\cong R_{Int,f}^{k}$ (2)
of weight n1n-1 and their dimensions, where $R_{Int,f}^{k}$ denotes a graded component of an interior module defined via jacobian rings (of Batyrev), (see [Bat93]). The isomorphism (2) identifies the mixed Hodge components of minimal weight with vector spaces defined by the lattice geometry of integral multiples of $\Delta$ (definition 2.5). In this article we find a representation of $R_{Int,f}^{k}$ as quotient vector space, which has not been regarded yet. This illustrates Griffiths concept of (infinitesimal) variation of Hodge structure for toric hypersurfaces: (Proposition 3.3)
Proposition 1.1.
Let $\Delta$ be an $n$-dimensional lattice polytope with $l^{*}(\Delta)>0$ and a given $f$. Given $\Gamma_{1},...,\Gamma_{n+1}\leq\Delta$ with $n_{\Gamma_{1}},...,n_{\Gamma_{n+1}}$ affine linear independent. Then
$R_{Int,f}^{k}=L^{*}(k\cdot\Delta)/U_{f,k}\quad k=1,...,n+1$
where $U_{f,k}$ denotes the vector space over $\mathbb{C}$ spanned by
$g_{\Gamma_{i}}(f)\cdot x^{v}\quad i=1,...,n+1,\quad v\in\operatorname{Int}((k-1)\cdot\Delta)\cap M$ (3)
$g_{\Gamma}(f)\cdot x^{v}\quad\Gamma\leq\Delta,\quad\quad\quad\quad\,\,\,\,v\in\operatorname{Int}((k-1)\cdot\Gamma)\cap M$ (4)
with certain Laurent polynomials $g_{\Gamma}(f)$. If $k=2$ these polynomials are linearly independent. Here $\operatorname{Int}$ denotes the relative interior and $L^{*}(k\cdot\Delta)$ the $\mathbb{C}$-vector space with basis $x^{m}$, where $m\in\operatorname{Int}(k\cdot\Delta)\cap M$.
The $g_{\Gamma}(f)$’s in fact are useful generators of the first component $J_{\Delta,f}^{1}$ of the jacobian ideal. These formulas heavily rely on the nondegeneracy of $f$ which could be translated in an algebraic independence between certain Laurent polynomials $F_{0},...,F_{n}$ (see theorem 2.4). A deficit is that the above generators of $U_{f,k}$ are in general not linearly independent for $k\geq 3$.
We then introduce the usual period map $\mathcal{P}_{B,f}$ with entries
$\mathcal{P}_{B,f}^{k}:f^{\prime}\mapsto F^{k}H^{n-1}(Y_{f^{\prime}},\mathbb{C})\subset H^{n-1}(Y_{f},\mathbb{C})$
with $\{F^{k}\}$ the Hodge filtration. This map is known to be holomorphic. The differential $d\mathcal{P}_{f}^{k}$ factors as follows
$T_{B,f}$$H^{1}(Y_{f},T_{Y_{f}})$$\operatorname{Hom}(H^{k}(Y_{f},\Omega_{Y_{f}}^{n-1-k}),H^{k+1}(Y_{f},\Omega_{Y_{f}}^{n-2-k}))$$\scriptstyle{d\mathcal{P}_{B,f}^{k}}$$\scriptstyle{\kappa_{f}}$$\scriptstyle{\Phi_{f}^{k}}$ (5)
Our idea is strainghtforward: Compute the kernel $\ker(d\mathcal{P}_{B,f}^{k})$ and compare it with $\ker(\kappa_{f})$, which we explicitly computed in the preceeding article [Gie22], to finally deduce on $\ker(\Phi_{f|{Im(\kappa_{f})}}^{k})$. The key of our results is the following theorem, which allows a computational approach (theorem 8.4)
Theorem 1.2.
Let $\Delta$ be an $n$-dimensional lattice polytope with $\operatorname{Int}(\Delta)\cap M\nsubseteq\textrm{ hyperplane}$. Then
$\ker(d\mathcal{P}_{B,f}^{k})\overset{\mod\operatorname{Lie}(T)}{=}$ $\Big\langle g_{\Gamma}(f)\cdot x^{w}\in L(\Delta)/\mathbb{C}\cdot f\quad\mid\quad\Gamma\leq\Delta\textrm{ a facet},$
$w+v\in\big(\operatorname{Int}((k+1)\cdot\Delta)\cup\operatorname{Int}((k+1)\cdot\Gamma)\big)\cap M,\quad$
$\forall\,v\in\operatorname{Int}(k\cdot\Delta)\cap M\Big\rangle.$
where $\langle\rangle$ denotes the span as $\mathbb{C}$-vector space.
This direct approach to the diagram of Griffiths (5) is not published yet and yields a both computational and much explicit perspective. Given the assumptions of the theorem we easily deduce
$\ker(d\mathcal{P}_{B,f}^{k})=\ker(d\mathcal{P}_{B,f}^{1}),\quad k\geq 1$
(Corollary 8.8). Let $g=g_{\Gamma}(f)\cdot x^{w}\in\ker(d\mathcal{P}_{f}^{1})$. Then we get the distinction
$\langle w,n_{\Gamma}\rangle=\left\{\begin{array}[]{lll}0,&&\Leftrightarrow g\equiv 0\\ -1,&&\Leftrightarrow g\in\ker(\kappa_{f})\\ \leq-2,&&\quad\,\,\textrm{exceptional cases}\end{array}\right$.} (6)
and $\langle w,n_{\Gamma^{\prime}}\rangle\geq 0$ for $\Gamma^{\prime}\neq\Gamma$ by Theorem (1.2) with $n_{\Gamma}$ the inner facet normal to $\Gamma$. The last case in the distinction causes $\Phi_{f|{Im\,\kappa_{f}}}$ to be not injective, which implies in particular that the infinitesimal Torelli theorem (ITT) fails for such hypersurfaces $Y_{f}$.
For smooth projective hypersurfaces $Y_{f}\subset\mathbb{P}^{n}$ of degree $\geq n+2$, where $n\geq 3$ the Kodaira-Spencer map $\kappa_{f}$ is surjective and we use this to get a new proof of the ITT for such nondegenerate hypersurfaces. We then finish with the following corollary
Corollary 1.3.
Given $n=3$ assume $\operatorname{Int}(\Delta)\cap M\nsubseteq\textrm{plane}$. Then
$\ker(\Phi_{f|{Im\,\kappa_{f}}})=\{0\},$
We say that $Y_{f}$ fulfills the ITT for $\mathcal{X}\rightarrow B$.
This corollary follows easily from above: Namely we construct a $3$-dimensional empty polytope $Q$ with $6$ vertices. $Q$ is the convex span of an empty triangle with vertex $(0,0,0)$ and this triangle dilated by $w$. Then we use a theorem of White on empty $3$-simplices to deduce $|Q\cap M|\geq 7$ under the assumption $\langle w,n_{\Gamma}\rangle\leq-2$, contradiction.
Our article relies on results from ([DK86]) and ([Bat93]). Let us list common methods with the authors V.I. Danilov, A.G. Khovanskii and V.V. Batyrev:
  • We reduce the calculation of the Hodge component $H^{n-k,k-1}(Y_{f},\mathbb{C})$ to the calculation of (2) (similarly to [DK86]), see section 5.
  • We follow [Bat93] to define the jacobian ring $R_{f}$ and the (interior) module $R_{Int,f}$ over $R_{f}$ (see section 2) and of course use the identification (2).
In section (7) we study $\Delta$ up to replacing by an integral lattice vector. In particular we effort $(0,...,0)\in\operatorname{Int}(\Delta)$ to build a smooth quotient of $B$ by the torus $T$.
Concerning the cokernel $\operatorname{coker}(\kappa_{f})$ the situation might get more complicated. In ([Mav03]) there were made some efforts for $Y_{f}$ a K3 hypersurface (so called non-polynomial deformations), realizing these deformations as deformations of complete intersections in a single big toric variety, but is seems unlikely that this representation is useful in order to compute the missing vector space $\ker(\Phi_{f|{\operatorname{coker}(\kappa_{f})}})$. In ([BrGr12]) there were made some efforts to combine the Hodge components and their computation to the generic Picard number of the hypersurface.
2 The Jacobian ring of Batyrev
In the sections 2 and 3 we do not restrict to the dimension $n=3$.
Definition 2.1.
Let $\Delta$ be an $n$-dimensional lattice polytope and $S_{\Delta}$ denote the subalgebra of $\mathbb{C}[x_{0},x_{1}^{\pm},...,x_{n}^{\pm}]$ spanned as $\mathbb{C}$-vector space by $1$ and all monomials
$x_{0}^{k}x_{1}^{m_{1}}...x_{n}^{m_{n}},\quad\textrm{where }k\in\mathbb{N}_{\geq 1},\quad m_{1},...,m_{n}\in\mathbb{Z},$
such that the rational point
$\frac{m}{k}:=\Big(\frac{m_{1}}{k},...,\frac{m_{n}}{k}\Big)$
belongs to $\Delta$. Denote the $k$-th graded piece of $S_{\Delta}$ by $S_{\Delta}^{k}$. The subalgebra $S_{\Delta}^{*}$ is defined in the same way except that we require that $m/k$ belongs to the interior of $\Delta$.
We identify lattice points $m\in M$ with monomials $x^{m}\in\mathbb{C}[x_{1}^{\pm 1},...,x_{n}^{\pm 1}]$.
Construction 2.2.
[Bat93]
Given a Laurent polynomial $f$ in the variables $x_{1},...,x_{n}$ with $n$-dimensional Newton polytope $\Delta$ let
$F(x_{0},x_{1},...,x_{n}):=x_{0}f(x_{1},...,x_{n})-1$
which is an equation for the complement $T\setminus Z_{f}\subset\overline{T}:=(\mathbb{C}^{*})^{n+1}$ of $Z_{f}=\{f=0\}\subset T$. Consider the logarithmic derivatives
$F_{i}(x_{0},x):=x_{i}\frac{\partial}{\partial x_{i}}F(x_{0},x)=x_{i}\frac{\partial(x_{0}f)}{\partial x_{i}}\quad 0\leq i\leq n.$
Definition 2.3.
[Bat93]
The graded ideal $J_{\Delta,f}$ within $S_{\Delta}$ generated by $F_{0},...,F_{n}$ is called the jacobian ideal and the quotient ring
$R_{f}:=S_{\Delta}/J_{\Delta,f}$
is called the Jacobian ring of Batyrev (associated to the Laurent polynomial $f$). We denote the $k$-th homogeneous component of $R_{f}$ by $R_{f}^{k}$.
Theorem 2.4.
[Bat93, Thm.4.8]
The jacobian ring $R_{f}$ is a graded ring. It is finite dimensional as $\mathbb{C}$-vector space if and only if $f$ is nondegenerate. In this case the dimensions of $R_{f}^{k}$ are independent of the polynomial $f$. $f$ is nondegenerate if and only if $F_{0},...,F_{n}$ are algebraically independent over $\mathbb{C}$.
Definition 2.5.
We denote the homogeneous ideal $R_{f}\cap S_{\Delta}^{*}$ of $R_{f}$ by $R_{Int,f}$ and its k-th homogeneous component by $R_{Int,f}^{k}$. We call $R_{Int,f}$ the interior $R_{f}$-module.
Remark 2.6.
Poincaré duality on $H^{n-1}(Y,\mathbb{C})$ restricts to an isomorphism
$(R_{Int,f}^{k})^{*}\cong R_{Int,f}^{n+1-k}$
by ([Bat93, Remark 9.5, Prop.9.7]).
3 Construction of the components $R_{Int,f}^{k}$
Remark 3.1.
By definition
$R_{Int,f}^{1}\cong L^{*}(\Delta)$
(independently of $f$). More interestingly
$R_{Int,f}^{k}\cong L^{*}(k\Delta)\big/\,\Big(J_{\Delta,f}^{k}\cap L^{*}(k\Delta)\Big),$ (7)
Since $J_{\Delta,f}^{k}$ is a graded ideal of $S_{\Delta}$ inductively
$J_{\Delta,f}^{k}=L((k-1)\Delta)\cdot J_{\Delta,f}^{1}$
Aim: Switching to different generators of $J_{\Delta,f}^{1}$ allows us to describe the relations in $R_{Int,f}^{k}$ more explicitly.
[…]
Figure 1: Illustration of the construction of $R_{Int,f}^{2}$ for a simplex $\Delta=\langle v_{1},v_{2},v_{3}\rangle$ in dimension $2$. The shaded regions do not belong to $R_{Int,f}^{2}$, there are 44 points left in $R_{Int,f}^{2}$.
Construction 3.2.
Let $\Delta$ be an $n$-dimensional lattice polytope with given
$f=\sum\limits_{m\in M\cap\Delta}a_{m}\cdot x^{m}.$
Given a facet $\Gamma=\Gamma_{i}\leq\Delta$, where $i\in\{1,...,r\}$, define
$g_{\Gamma}(f):=\sum\limits_{m\in M\cap\Delta}a_{m}\cdot(\langle n_{i},m\rangle+b_{i})\cdot x_{0}x^{m}\underbrace{=}_{(\textrm{Trafo.})}+b_{i}F_{0}+\sum\limits_{j=1}^{n}(n_{i})_{j}F_{j}\in J_{\Delta,f}^{1},$
where $n_{i}=((n_{\Gamma})_{1},...,(n_{\Gamma})_{n})$ and $b_{i}=-\operatorname{Min}_{\Delta}(n_{i})$. The first representation implies that
$\operatorname{Supp}(g_{\Gamma}(f))\subset\big(\Delta\cap M\big)\setminus\big(\Gamma\cap M\big).$
In case $\Delta$ is an $n$-simplex and
$\operatorname{Supp}(f)=\textrm{Vert}(\Delta)$
then $g_{\Gamma_{i}}(f)\approx x^{v_{i}}$ where $v_{i}$ is the vertex opposite to $\Gamma_{i}$ ($\approx$ means up multiplication with a nonzero scalar). Conversely the matrix
$\begin{pmatrix}b_{1}&(n_{\Gamma_{1}})_{1}&...&(n_{\Gamma_{1}})_{n}\\ \vdots&\vdots\\ b_{r}&(n_{\Gamma_{r}})_{1}&...&(n_{\Gamma_{r}})_{n}\end{pmatrix}$ (8)
has rank $n+1$ since $n_{\Gamma_{1}},...,n_{\Gamma_{r}}$ span $N_{\mathbb{R}}$ and $(b_{1},...,b_{r})\neq(0,...,0)$. Thus $F_{0},...,F_{n}$ are linear combinations of $g_{\Gamma_{1}}(f),...,g_{\Gamma_{r}}(f)$ and we get new generators
$J_{\Delta,f}=(g_{\Gamma_{1}}(f),...,g_{\Gamma_{r}}(f)).$ (9)
Proposition 3.3.
Let $\Delta$ be an $n$-dimensional lattice polytope with $l^{*}(\Delta)>0$ and a given $f$. Given $\Gamma_{1},...,\Gamma_{n+1}\leq\Delta$ with $n_{\Gamma_{1}},...,n_{\Gamma_{n+1}}$ affine linear independent. Then
$R_{Int,f}^{k}=L^{*}(k\cdot\Delta)/U_{f,k}\quad k=1,...,n+1$
where $U_{f,k}$ denotes the vector space over $\mathbb{C}$ spanned by
$g_{\Gamma_{i}}(f)\cdot x^{v}\quad i=1,...,n+1,\quad v\in\operatorname{Int}((k-1)\cdot\Delta)\cap M$ (10)
$g_{\Gamma}(f)\cdot x^{v}\quad\Gamma\leq\Delta,\quad\quad\quad\quad\,\,\,\,v\in\operatorname{Int}((k-1)\cdot\Gamma)\cap M$ (11)
If $k=2$ these polynomials are linearly independent.
Proof.
The inclusion
$U_{f,k}\subseteq\Big(J_{\Delta,f}^{k}\cap L^{*}(k\Delta)\Big).$
is a consequence of the definition of $J_{\Delta,f}$.
To show: All relations
$h\cdot g\in L^{*}(k\cdot\Delta),\quad\quad\quad h\in J_{\Delta,f}^{1},\quad g\in L((k-1)\cdot\Delta)$ (12)
are of type (10) or (11). First if $\operatorname{Supp}(g)\subseteq\operatorname{Int}((k-1)\cdot\Delta)\cap M$ then the equation 12 is a linear combination of the relations (10) and without restriction
$\operatorname{Supp}(g)\subset\operatorname{Bound}((k-1)\cdot\Delta)\cap M.$
Write
$h=\sum\limits_{\Gamma}c_{\Gamma}\cdot g_{\Gamma}(f)$
and take $F\leq\Delta$ with
$\operatorname{Supp}(g)\cap\Big((k-1)\cdot F\cap M\Big)\neq\emptyset,$
(this assumption makes sense since $\operatorname{Supp}(g)\neq\emptyset$). Then
$\operatorname{Supp}(h)\cap(F\cap M)=\emptyset$
by (12). In case $c_{\Gamma}=0$ except for $\Gamma=F$, then
$\operatorname{Supp}(g)\subset((k-1)\cdot F)$
and we get a relation of type (11). Restricting $h$ to $F\cap M$ we get
$h_{|{F\cap M}}=\sum\limits_{\Gamma}c_{\Gamma}\cdot\big(-b_{\Gamma}F_{0|{\Gamma_{1}}}+\sum\limits_{i=1}^{n}(n_{\Gamma})_{i}F_{i|{\Gamma_{1}}}\big)\in\langle F_{0|{\Gamma_{1}}},...,F_{n|{\Gamma_{1}}}\rangle.$
Expanding and restricting to $F$ this means
$\sum\limits_{\Gamma}c_{\Gamma}\sum\limits_{m\in F\cap M}a_{m}\big(\langle n_{\Gamma},m\rangle-b_{\Gamma}\big)x_{0}x^{m}=0.$ (13)
The left hand side in (13) equals
$\sum\limits_{\Gamma}c_{\Gamma}\cdot\big(-b_{\Gamma}F_{0|{\Gamma_{1}}}+\sum\limits_{i=1}^{n}(n_{\Gamma})_{i}F_{i|{\Gamma_{1}}}\big)\in\langle F_{0|{\Gamma_{1}}},...,F_{n|{\Gamma_{1}}}\rangle.$
The $F_{i}$ are algebraically independent over $\mathbb{C}$ (Theorem 2.4), since $f$ is nondegenerate with respect to $\Delta$. Besides $f_{|{\Gamma_{1}}}$ remains nondegenerate with respect to $\Gamma_{1}$ (by the definition of nondegeneracy) and it follows that there is (up to scaling) only one relation between $F_{0|{\Gamma_{1}}},...,F_{n|{\Gamma_{1}}}$, the one of the second type.
For $k=2$ we know the dimension ([Bat93, Thm.9.8])
$\dim\,R_{Int,f}^{2}=l^{*}(2\cdot\Delta)-(n+1)\cdot l^{*}(\Delta)-\sum\limits_{\Gamma}l^{*}(\Gamma),$
where sum ranges over all facets $\Gamma$ of $\Delta$. Thus the last statement follows by comparing dimensions. ∎
Example 3.4.
Let $\Delta$ be a simplex with vertices $v_{0},...,v_{n}$ and assume that $f$ has support on the vertices of $\Delta$. Then as already noted
$g_{\Gamma_{i}}(f)=x^{v_{i}},$
where $\Gamma_{i}$ denotes the facet opposite to $v_{i}$. In this case we get a monomial basis for example of $R_{Int,f}^{2}$ by taking the quotient of $L^{*}(2\cdot\Delta)$ by the span of
$x^{v_{i}+v}\quad i=0,...,n,\quad v\in\big(\operatorname{Int}(\Delta)\cup\operatorname{Int}(\Gamma_{i})\big)\cap M.$
Remark 3.5.
For $k>2$ the polynomials in the Proposition will not be linear independent over $\mathbb{C}$ since we have the trivial relations
$g_{\Gamma_{i}}(f)\cdot g_{\Gamma_{j}}(f)\cdot x^{v}-g_{\Gamma_{j}}(f)\cdot g_{\Gamma_{i}}(f)\cdot x^{v}$
for $v\in L^{*}((k-2)\Delta)$.
4 Hodge and mixed Hodge structures
In this section we give some theoretical background on Hodge and mixed Hodge structures.
Definition 4.1.
Let $H$ be a finite dimensional $\mathbb{Q}$-vector space and $H_{\mathbb{C}}:=H\otimes\mathbb{C}$. A Hodge structure of weight $k$ on $H$ is a decomposition
$H_{\mathbb{C}}=\bigoplus\limits_{p+q=k}H^{p,q},$
such that $H^{p,q}=\overline{H^{q,p}}$, where the bar denotes complex conjugation on $H_{\mathbb{C}}$.
Remark 4.2.
To put a Hodge structure of weight $k$ on $H$ is equivalent to the existence of a descending filtration
$H_{\mathbb{C}}=F^{0}\supset F^{1}\supset...\supset F^{k+1}=0$
with $H_{\mathbb{C}}=F^{p}\cap\overline{F^{k+1-p}}$. This correspondence is given by
$F^{i}=\bigoplus\limits_{p\geq i}H^{p,k-p},\quad\textrm{and conversely by}\quad H^{p,q}=F^{p}\cap\overline{F^{q}}.$
Definition 4.3.
Let $H$ be a finite dimensional $\mathbb{Q}$-vector space, $H_{\mathbb{C}}:=H\otimes\mathbb{C}$. A mixed Hodge structure (for short: MHS) of weight $k$ on $H$ consists of an ascending (weight) filtration
$0=W_{k-1}\subset W_{k}\subset...\subset W_{2k}=H$
and a descending (Hodge) filtration
$H_{\mathbb{C}}=F^{0}\supset F^{1}\supset...\supset F^{k+1}=0$
such that the Hodge filtration induces a Hodge structure of weight $r$ on $W_{r}/W_{r-1}$.
We denote the successive quotients $F^{i+1}/F^{i}$ by $\operatorname{Gr}_{F}^{i}$ and $W_{r}/W_{r-1}$ by $\operatorname{Gr}_{W}^{r}$. A Hodge structure of weight $k$ on $H$ defines also a mixed Hodge structure on $H$ by setting $W_{k}:=H$. These abstract definitions are motivated by the following fact:
Theorem 4.4.
([Del75], see also [Dan78, §14])
For $Y$ an $n$-dimensional quasismooth hypersurface in a simplicial projective toric variety $\mathbb{P}$ the cohomology groups $H^{k}(Y,\mathbb{Q})$ are equipped with a Hodge structure of weight $k$ and we have
$H^{p,k-p}(Y):=H^{p,k-p}\cong H^{k-p}(Y,\Omega_{Y}^{p}).$
For $\Delta$ an $n$-dimensional lattice polytope, $f\in U_{reg}(\Delta)$ and $Z_{f}=\{f=0\}\subset T$ the cohomology groups
$H^{k}(Z_{f},\mathbb{Q})$
carry a mixed Hodge structure of weight $k$.
If $H=H^{k}(Y,\mathbb{Q})$ we write $F^{i}H^{k}(Y,\mathbb{C})$ and $W_{j}H^{k}(Y,\mathbb{Q})$ for $F^{i}$ and $W_{j}$.
Definition 4.5.
The dimensions
$h^{p,k-p}(Y):=\dim\,H^{p,k-p}(Y)$
are called the Hodge numbers of $Y$. The spaces $H^{p,k-p}(Y)$ and
$H^{i,j+k-i}H^{k}(Z_{f},\mathbb{C}):=\operatorname{Gr}_{F}^{i}\operatorname{Gr}_{W}^{j+k}H^{k}(Z_{f},\mathbb{C}),$
are called the Hodge components of $Y$ and $Z_{f}$.
Definition 4.6.
Let $H,H^{\prime}$ be $\mathbb{Q}$-vector spaces carrying mixed Hodge structures of weight $k$ and $k+2\cdot l$ given by filtrations
$(F^{i},W_{j})\quad\textrm{and}\quad((F^{\prime})^{i},W^{\prime}_{j}).$
A $\mathbb{Q}$-linear homomorphism $\phi:H\rightarrow H^{\prime}$ is called a morphism of mixed Hodge structures of type $(l,l)$ if
$\phi(W_{r})\subset W^{\prime}_{r+2l},\quad\phi(F^{p})\subset(F^{\prime})^{p+l}.$
Remark 4.7.
For $Y$ a smooth minimal model of $Z_{f}$ the inclusion $j:Z_{f}\rightarrow Y$ induces a pullback homomorphism
$j^{*}:H^{k}(Y,\mathbb{C})\rightarrow H^{k}(Z_{f},\mathbb{C}).$
If we equip $H^{k}(Y,\mathbb{Q})$ with the trivial weight filtration
$W_{k}H^{k}(Y,\mathbb{Q})=H^{k}(Y,\mathbb{Q}),$
then $j^{*}$ gets a morphism of MHS of type $(0,0)$ ([Voi02, Ch.7]). This means that the restrictions of cohomology classes of $Y$ to $Z_{f}$ lie in the minimal weight subspace $W_{k}H^{k}(Z_{f},\mathbb{C})$. Cohomology classes in $W_{k+j}H^{k}(Z_{f},\mathbb{C})$ for $j>0$ vanish on the comactification $Y$ of $Z_{f}$.
5 The Hodge components of $Z_{f}$ and $Y_{f}$
Definition 5.1.
We define the primitive cohomology of $Z_{f}$ as follows:
$PH^{n-1}(Z_{f},\mathbb{C}):=\operatorname{coker}(H^{n-1}(T,\mathbb{C})\rightarrow H^{n-1}(Z_{f},\mathbb{C})).$
Remark 5.2.
There are induced filtrations $F^{i}$ and $W_{j}$ on $PH^{n-1}(Z_{f},\mathbb{C})$. The torus $T$ is homotopy equivalent to a topological torus, thus
$H^{i}(T,\mathbb{C})\cong\wedge^{i}M_{\mathbb{C}}.$
In particular $h^{i}(T,\mathbb{C})=\binom{n}{i}$. By the Lefschetz theorem for hypersurfaces in tori ([DK86, Remark 3.10]) the restriction homomorphism
$H^{i}(T,\mathbb{C})\rightarrow H^{i}(Z_{f},\mathbb{C})$
is an isomorphism for i<n1i<n-1 and injective for i=n1i=n-1.
Remark 5.3.
We have an isomorphism of $\mathbb{C}$ vector spaces ([Bat93, Thm.6.9, Cor.6.10])
$PH^{n-1}(Z_{f},\mathbb{C})\cong R_{f}.$
This allows us to transport the Hodge and the weight filtration to $R_{f}$ (see [Bat93, Thm.6.9, Thm.8.2]): The Hogde filtration is given by the reverse grading on $R_{f}$ and the weight filtration on $R_{f}^{k}$ is induced by the subdivision of $k\cdot\Delta$ into $j$-dimensional faces for j=0,,nj=0,...,n. In this thesis we just need the following result:
Theorem 5.4.
([Bat93, Prop.9.2])
The Hodge components
$H^{p,n-1-p}H^{n-1}(Z_{f},\mathbb{C}):=Gr_{F}^{p}Gr_{W}^{n-1}H^{n-1}(Z_{f},\mathbb{C})$
are computed as follows:
$H^{p,n-1-p}H^{n-1}(Z_{f},\mathbb{C})\cong R_{Int,f}^{n-p}.$
Remark 5.5.
(see [Bat22], [Gie22, section 2])
Let $\Delta$ be an $n$-dimensional lattice polytope with $F(\Delta)\neq\emptyset$. The the restrictions of prime toric boundary divisors of $\mathbb{P}_{C(\Delta)}$ and the restrictions of toric strata which are exceptional for $\rho\circ\pi:\mathbb{P}\rightarrow\mathbb{P}_{C(\Delta)}$ compactify $Z_{f}$ to $Y_{f}$. We denote the divisorial components of all these divisors by $F_{1},...,F_{r}$ and the sum of the components of codimension $\geq 2$ in $Y_{f}$ by $S$.
Let $Z_{f}^{*}:=Y_{f}\setminus D$, where
$D:=F_{1}+...+F_{r}.$
There is the following Gysin exact sequence (compare also [DK86, Proof of Thm.3.7]):
$...\rightarrow H^{n-2}(Y_{f},\mathbb{C})$ $\rightarrow H^{n-2}(Z_{f}^{*},\mathbb{C})\overset{r}{\rightarrow}\bigoplus\limits_{i=1}^{r+s}H^{n-3}(F_{i},\mathbb{C})\overset{k_{*}}{\rightarrow}H^{n-1}(Y_{f},\mathbb{C})$
$\overset{j^{*}}{\rightarrow}H^{n-1}(Z_{f}^{*},\mathbb{C})\overset{r}{\rightarrow}\bigoplus\limits_{i=1}^{r+s}H^{n-2}(F_{i},\mathbb{C})\rightarrow.$...
The homomorphism $j^{*}$ comes from the inclusion $j:Z_{f}\xhookrightarrow{}Y_{f}$, $r$ is the residue map and $k_{*}$ the so called Gysin map. If we equip $H^{n-1}(Y_{f},\mathbb{C})$ with the trivial weight filtration then the map $j^{*}$ is a morphism of MHS of type $(0,0)$ (Remark 4.7), the Gysin map is of Hodge type $(1,1)$, whereas $r$ is of Hodge-type (1,1)(-1,-1) (see [Voi02, Ch.7]).
Restricting to $\operatorname{Gr}_{W}^{n-1}$ we get
$\operatorname{Gr}_{W}^{n-1}H^{n-2}(Y_{f},\mathbb{C})=0,\quad\operatorname{Gr}_{W}^{n-1}H^{n-2}(F_{i},\mathbb{C})=0.$
We arrive at
$0$ $\rightarrow Gr_{W}^{n-1}H^{n-2}(Z_{f}^{*},\mathbb{C})\overset{r}{\rightarrow}\bigoplus\limits_{i=1}^{r}\operatorname{Gr}_{W}^{n-3}H^{n-3}(F_{i},\mathbb{C})\overset{k_{*}}{\rightarrow}H^{n-1}(Y_{f},\mathbb{C})$ (14)
$\overset{j^{*}}{\rightarrow}Gr_{W}^{n-1}H^{n-1}(Z_{f}^{*},\mathbb{C})\rightarrow 0.$
Comparing $Z_{f}^{*}$ with $Z_{f}$ we immediately get
$H^{n-1}(Z_{f},\mathbb{C})\cong H^{n-1}(Z_{f}^{*},\mathbb{C}).$
Theorem 5.6.
Let $\Delta$ be an $n$-dimensional lattice polytope with $F(\Delta)\neq\emptyset$. Let $Y_{f}$ be a minimal model of $Z_{f}$. Then
$H^{p}(Y_{f},\Omega_{Y_{f}}^{n-1-p})\approx R_{Int,f}^{p+1},$ (15)
where $\approx$ means up to some cohomology classes that arise as restrictions from cohomology classes from strata of $\mathbb{P}$ and which stay constant on the whole family $\mathcal{X}\overset{pr_{2}}{\rightarrow}B$.
6 Properties of the period map and its differential
Definition 6.1.
([Voi02, Thm.9.3, Ch.10.1.2, Ch.10.1.3])
Let $\Delta$ be an $n$-dimensional lattice polytope with $l^{*}(\Delta)>0$, let $f\in B$ and assume $Y_{f}$ is smooth. The period map $\mathcal{P}_{B,f}$ for the n1n-1-th cohomology is defined by
$\mathcal{P}_{B,f}=\bigoplus\limits_{k=1}^{n-1}\mathcal{P}_{B,f}^{k}:$ $B\rightarrow\Gamma\setminus D$
$f^{\prime}\mapsto(F^{n-1}H^{n-1}(Y_{f^{\prime}},\mathbb{C}),...,F^{1}H^{n-1}(Y_{f^{\prime}},\mathbb{C}))$
where
$H^{n-k,k-1}(Y_{f^{\prime}})\subset H^{n-1}(Y_{f^{\prime}},\mathbb{C})\cong H^{n-1}(Y_{f},\mathbb{C})$
Here
$D:\textrm{ period domain (a quasiprojective variety) }$
$\Gamma:\textrm{The monodromy group }\pi_{1}(B,f)$
We refer to ([Voi02, Ch.10], [Voi03, Ch.3]) for details.
Construction 6.2.
(Result of Griffiths) ([Voi02, Thm.10.21])
The differential $d\mathcal{P}_{B,f}$ of $\mathcal{P}_{B,f}$ fits into a diagram
$T_{B,f}$$H^{1}(Y_{f},T_{Y_{f}})$$\bigoplus\limits_{k=0}^{n-2}\operatorname{Hom}(H^{k}(Y_{f},\Omega_{Y_{f}}^{n-1-k}),H^{k+1}(Y_{f},\Omega_{Y_{f}}^{n-2-k}))$$\scriptstyle{d\mathcal{P}_{B,f}}$$\scriptstyle{\kappa_{f}}$$\scriptstyle{\Phi_{f}}$ (16)
since $Y_{f}$ is smooth. $\Phi_{f}$ is the homomorphism between cohomology groups induced by cup product and the contraction
$T_{Y_{f}}\times\Omega_{Y_{f}}^{n-1-k}\rightarrow\Omega_{Y_{f}}^{n-2-k}.$
This diagram is important since it connects the Hodge-theoretic homomorphism $d\mathcal{P}_{B,f}$ with the Kodaira-Spencer map $\kappa_{f}$.
Remark 6.3.
Starting with a smooth proper deformation $\mathcal{Y}\rightarrow S$ of $Y_{f}$ with $S$ smooth, we define a Kodaira Spencer map $\kappa_{S,f}$ and a period map $\mathcal{P}_{S,f}$ just as in the Definitions above. The result of Griffiths remains valid: $d\mathcal{P}_{S,f}$ factors through $\kappa_{S,f}$ and $\Phi_{f}$, that is $\Phi_{f}$ is universal.
Definition 6.4.
The infinitesimal Torelli Theorem (short: ITT) for $Y_{f}$ asks if $\Phi_{f}$ is injective. The infinitesimal Torelli Theorem for $Y_{f}$ in $\mathcal{X}\overset{pr_{2}}{\rightarrow}B$ asks if $\Phi_{f|{Im\,\kappa_{f}}}$ is injective. The infinitesimal Torelli Theorem for $\mathcal{X}\overset{pr_{2}}{\rightarrow}B$ asks if $\Phi_{f|{Im\,\kappa_{f}}}$ is injective for $f\in B$.
If $\kappa_{f}$ is surjective of course the first and the third definition coincide. Given the identifications (5.4) choose a reference point $f\in B$ and define a mixed period map $\phi_{f}$ from $B$ into a mixed period domain $D_{mix}$ (which has to be defined) by
$f^{\prime}\overset{\phi_{f}}{\mapsto}(R_{Int,f^{\prime}}^{1},...,R_{Int,f^{\prime}}^{n}).$
Apparently $\phi_{f}$ just depends on the affine part $Z_{f}$ and not on the particular compactification $Y_{f}$.
Lemma 6.5.
$d\phi_{f}^{k}$ has image in $\operatorname{Hom}(R_{Int,f}^{k},R_{Int,f}^{k+1})$ and $d\mathcal{P}_{B,f}^{k}$ factors as follows
$T_{B,f}$$\operatorname{Hom}(R_{Int,f}^{k},R_{Int,f}^{k+1})$$\operatorname{Hom}(H^{k-1}(Y_{f},\Omega_{Y_{f}}^{n-k}),H^{k}(Y_{f},\Omega_{Y_{f}}^{n-k+1}))$$\scriptstyle{d\mathcal{P}_{B,f}^{k}}$$\scriptstyle{d\phi_{f}^{k}}$ (17)
where the vertical map denotes the inclusion.
7 Smooth and stable points of $\mathcal{M}(\Delta)$
Given an $n$-dimensional lattice polytope $\Delta$ and some polynomial $f$, the torus $T$ acts on $U_{reg}(\Delta)$
$(t_{1},...,t_{n}).f(x_{1},...,x_{n})=f(t_{1}x_{1},...,t_{n}x_{n}),\quad(t_{1},...,t_{n})\in T.$
$T\cdot f$ denotes the orbit of $f$ under $T$.
Let $\mathcal{M}(\Delta):=B/T$ be the quotient of $B$ by $T$. We omit equivalence classes and write $f\in\mathcal{M}(\Delta)$.
Definition 7.1.
([MuFo82, Def.1.7])
Let $v\in\mathbb{C}^{l(\Delta)+1}$ and $x=[v]$. The point $x$ is stable if the orbit $T\cdot v$ is closed and of dimension $\dim\,\Delta=n$. The second condition is equivalent to the condition
$\operatorname{Stab}_{T}(v)<\infty,$
where
$\operatorname{Stab}_{T}(v):=\{t\in T|\,t.v=v\}$
denotes the stabilizer of $v$ w.r.t. the action of $T$.
Let $r:=l(\Delta)-1$. The set $(\mathbb{P}^{r})^{s}$ of stable points of $\mathbb{P}^{r}$ is Zariski open in $\mathbb{P}^{r}$ ([MuFo82, §4]), but might be empty. Let
$\Delta_{r}:=\langle e_{m}\,|\quad m\in M\cap\Delta\rangle$
denote the $r$-dimensional standard simplex embedded into an affine hyperplane in $\mathbb{R}^{l(\Delta)}$. Then there is a map $\pi:\Delta_{r}\rightarrow\Delta$ given by
$\sum\limits_{m\in M\cap\Delta}\lambda_{m}\cdot e_{m}\mapsto\sum\limits_{m\in M\cap\Delta}\lambda_{m}\cdot m\quad\textrm{for }\sum\limits_{m\in M\cap\Delta}\lambda_{m}=1.$
To $\Delta_{r}$ is associated the toric variety $\mathbb{P}^{r}$ and given
$a:=(a_{m})_{m\in M\cap\Delta}\in\mathbb{P}^{r}$
there is a natural $(\mathbb{C}^{*})^{r}$ orbit through a of some dimension $k\in\{0,...,r\}$. We denote the $k$-dimensional face of $\Delta_{r}$ corresponding to this orbit by $\Gamma(a)$.
Proposition 7.2.
([KSZ, Prop.3.5])
$(\mathbb{P}^{r})^{s}\neq\emptyset$ if and only if
$(0,...,0)\in\operatorname{Int}(\Delta)\cap M.$
More precisely $a\in(\mathbb{P}^{r})^{s}$ is stable if and only if $\pi(\Gamma(a))$ has full dimension $n$ and contains (0,,0)(0,...,0) in its interior.
Corollary 7.3.
If $(0,...,0)\in\operatorname{Int}(\Delta)\cap M$ and $a:=(a_{m})_{m\in M\cap\Delta}\in\mathbb{R}^{l(\Delta)}$ is such that $f$ lies in $U_{reg}(\Delta)$, then $a$ is stable.
Proof.
For $m$ a vertex of $\Delta$ we have $a_{m}\neq 0$ and thus $\pi(\Gamma(a))\supset\operatorname{Int}(\Delta)$, in particular $(0,...,0)\in\pi(\Gamma(a))$ and $\pi(\Gamma(a))$ is full-dimensional. ∎
Corollary 7.4.
If $(0,...,0)\in\operatorname{Int}(\Delta)$ then the quotient $\mathcal{M}(\Delta)$ is smooth.
Proof.
By definition $L(\Delta)\setminus U_{reg}(\Delta)=\{E_{A}=0\}$, where $A:=\Delta\cap M$. In effect $U_{reg}(\Delta)$ is affine.
The orbit $T\cdot f$ is closed and $n$-dimensional since all $f\in\mathcal{M}(\Delta)$ are stable w.r.t. the action of $T$ on $\mathbb{P}^{l(\Delta)-1}$. Applying Luna’s slice Theorem ([Dre12, Prop.5.7]) to the affine variety $U_{reg}(\Delta)$ and the projection
$U_{reg}(\Delta)\rightarrow\mathcal{M}(\Delta)$
gives us that $\mathcal{M}(\Delta)$ is smooth at $f$ if $\operatorname{Stab}_{T\times\mathbb{C}^{*}}(f)$ contains just the neutral element (1,,1)(1,...,1).
By definition
$\operatorname{Stab}_{T}(f)=\{(t,1)=(t_{1},...,t_{n},1)\in T\times\mathbb{C}^{*}\,|\quad t^{m}a_{m}=a_{m}\quad\forall m\,\textit{ with }a_{m}\neq 0\}.$
Thus
$\operatorname{Stab}_{T}(f)\subset\{t\in T\,|\,t^{v_{i}}=1\quad\forall\textrm{ vertices }v_{1},...,v_{k}\}\times\{1\}^{.}$
Consider $n$ vertices $v_{1},...,v_{n}$ which span $M_{\mathbb{R}}$ and apply an unimodular transformation $U:M\rightarrow M$ such that
$U(v_{1}),...,U(v_{n-1})\in\{(m_{1},...,m_{n})\in M|\,m_{n}=0\}.$
Replace $v_{i}$ by $U(v_{i})$ and note that for
$(t_{1},...,t_{n})\in\operatorname{Stab}_{T}(f)$
the entry $t_{n}$ is uniquely determined by $t_{1},...,t_{n-1}$ and the relation $1=t^{v_{n}}$. Further if $(t_{1},...,t_{n-1})=(1,...,1)$ then also $t_{n}=1$. By this we have reduced the assertion $\operatorname{Stab}_{T}(f)=\{(1,...,1)\}$ to the lower-dimensional problem that the only solution of
$1=t^{v_{1}}=...=t^{v_{n-1}}$
is t=(1,,1)t=(1,...,1). We continue inductively.
Remark 7.5.
Given $(0,...,0)\in\operatorname{Int}(\Delta)\cap M$ and $f\in\mathcal{M}(\Delta)$ then $f$ is smooth and stable. Let $T\rightarrow B$ denote the map
$(t_{1},...,t_{n})\mapsto f(t_{1}x_{1},...,t_{n}x_{n}).$
Then we get
$T_{\mathcal{M}(\Delta),f}$ $\cong T_{B,f}/T_{T\cdot f,f}$
$\cong L(\Delta)/\langle f,x_{1}\frac{\partial f}{\partial x_{1}},...,x_{n}\frac{\partial f}{\partial x_{n}}\rangle\cong R_{f}^{1}$
where the first isomorphism follows from ([Bat93, Cor.11.3]).
Remark 7.6.
Given $f\in\mathcal{M}(\Delta)$ the differential $d\phi_{f}$ and $\kappa_{f}$ factor through $T_{\mathcal{M}(\Delta),f}$. The differential $d\phi_{f}$ is simply induced by the addition of lattice points (see [Bat93, Prop.11.8])
[…]
8 Explicit description of $d\phi_{f}$
The following elementary Remark follows also from diagram 16.
Remark 8.1.
(Elementary proof that $\ker(\kappa_{f})\subset\ker(d\phi_{f}^{1}$))
We use the notation from ([Gie22]): Let $\alpha\in R(N,\Sigma_{\Delta})$. Then there is $\Gamma_{-\alpha}\leq\Delta$ such that if $v\in\operatorname{Int}(\Delta)\cap M$ then by definition of the roots
$v-\alpha\in\operatorname{Int}(\Delta)\cap M\quad\textrm{or}\quad v-\alpha\in\operatorname{Int}(\Gamma_{-\alpha})\cap M.$
We get
$g_{\Gamma_{-\alpha}}(f)\cdot x^{-\alpha}=w_{-\alpha}(f).$
and
$w_{-\alpha}(f)\cdot x^{v}=g_{\Gamma_{-\alpha}}(f)\cdot x^{v-\alpha}\in L^{*}(2\Delta).$
Since $g_{\Gamma_{-\alpha}}(f)\cdot x^{v-\alpha}\in J_{\Delta,f}^{2}$ we verified explicitly that $w_{-\alpha}(f)\in\ker(d\phi_{f}^{1})$.
The kernel $\ker(d\phi_{f})$:
Remark 8.2.
Assume $(0,...,0)\in\operatorname{Int}(\Delta)\cap M$. First by definition
$\ker(d\phi_{f}^{k})=\{h\in L(\Delta)\mid\,\forall v\in\operatorname{Int}(k\cdot\Delta)\cap M:\,h\cdot x^{v}\in J_{\Delta,f}^{k+1}\}.$ (19)
and by Proposition 3.3
$L^{*}((k+1)\cdot\Delta)\cap J_{\Delta,f}^{k+1}=$ $\langle g_{\Gamma}(f)\cdot x^{w}\mid\Gamma\leq\Delta,$
$w\in\operatorname{Int}(k\cdot\Delta)\cap M\textrm{ or }w\in\operatorname{Int}(k\cdot\Gamma)\cap M\rangle.$
In general given $h\in\ker(d\phi_{f}^{k})$
$h\cdot x^{v}=\sum\limits_{\Gamma}h_{\Gamma,v}\cdot g_{\Gamma}(f)$ (20)
for $v\in\operatorname{Int}(k\cdot\Delta)\cap M$ and $\operatorname{Supp}(h_{\Gamma,v})\subset(\operatorname{Int}(k\cdot\Delta)\cup\operatorname{Int}(k\cdot\Gamma))\cap M$.
Remark 8.3.
$h_{\Gamma,v}$ is completely determined by $h_{\Gamma,(0,0,0)}$: Given $h\in\ker(d\phi_{f}^{k})$ and $v\in\operatorname{Int}(k\cdot\Delta)\cap M$ there are $h_{\Gamma,v}\in\big(\operatorname{Int}(k\cdot\Delta)\cup\operatorname{Int}(k\cdot\Gamma)\big)\cap M$ such that
$h=\sum\limits_{\Gamma}h_{\Gamma,v}\cdot g_{\Gamma}(f)\cdot x^{-v}$
Setting $h_{\Gamma}:=h_{\Gamma,(0,0,0)}$ the relation (take $R$ be a set with $g_{\Gamma}(f)\cdot x^{v}$ a basis (linear independent) of $U_{f,k}$ for $(\Gamma,v)\in R$)
$\sum\limits_{(\Gamma,v)\in R}h_{\Gamma}\cdot g_{\Gamma}(f)\cdot x^{v}=h\cdot x^{v}=\sum\limits_{(\Gamma,v)\in R}h_{\Gamma,v}\cdot g_{\Gamma}(f)$
implies $h_{\Gamma,v}=h_{\Gamma}\cdot x^{v}$ by linear independence of the $g_{\Gamma}(f)$’s.
Theorem 8.4.
Let $\Delta$ be an $n$-dimensional lattice polytope with $\operatorname{Int}(\Delta)\cap M\nsubseteq\textrm{ hyperplane}$. Then
$\ker(d\phi_{f}^{k})=$ $\Big\langle g_{\Gamma}(f)\cdot x^{w}\in R_{f}^{1}\mid\Gamma\leq\Delta\textrm{ a facet},$
$w+v\in\big(\operatorname{Int}((k+1)\cdot\Delta)\cup\operatorname{Int}((k+1)\cdot\Gamma)\big)\cap M,\quad$
$\forall\,v\in\operatorname{Int}(k\cdot\Delta)\cap M\Big\rangle.$
The inclusion $\supseteq$ is obvious. The problem with the opposite inclusion is that it might happen that
$\sum\limits_{\Gamma}h_{\Gamma}\cdot g_{\Gamma}(f)\in L(\Delta).$ (21)
but $h_{\Gamma}\cdot g_{\Gamma}(f)\notin L(\Delta)$ for several $\Gamma$’s. We start with two Lemmas: Suppose given an element of the kernel as in (8.8). Let
$H_{n,l}:=\{x\in M_{\mathbb{R}}\mid\langle x,n\rangle=l\}\quad n\in N,\quad l\in\mathbb{Z}.$
Lemma 8.5.
Let $R_{\Gamma}:=(\operatorname{Int}(\Delta)\cup\operatorname{Int}(\Gamma))\cap M$.
$\operatorname{Supp}(h_{\Gamma})\subset\operatorname{Cone}(\operatorname{Int}(\Gamma)\cap M)\cap H_{n_{\Gamma},-l_{\Gamma}}\cap M,$ (22)
where $l_{\Gamma}$ denotes the smallest natural number $\geq 1$ with
$\emptyset\neq H_{n_{\Gamma},-l_{\Gamma}}\cap R_{\Gamma}$
(if $l_{\Gamma}$ does not exist, then $\operatorname{Supp}(h_{\Gamma})=\emptyset$).
Proof.
If
$(0,0,0)\neq m\in\operatorname{Supp}(h_{\Gamma})$
then some multiple $r\cdot m$, $r\in\mathbb{N}_{\geq 1}$, lies in $\operatorname{Int}(\Gamma)\cap M$ by (20). Thus $m\in\operatorname{Cone}(\operatorname{Int}(\Gamma)\cap M)$. If there were $m^{\prime}\in\operatorname{Int}(\Delta)\cap M$ with
$0>\langle m^{\prime},n_{\Gamma}\rangle>\langle m,n_{\Gamma}\rangle$
then
$(r-1)\cdot m+m^{\prime}\in\operatorname{Int}(k\cdot\Delta)\cap M$
$\Rightarrow\langle r\cdot m+m^{\prime},n_{\Gamma}\rangle=\Big(\underbrace{(r-1)\cdot m+m^{\prime}}_{\in\operatorname{Int}(\Delta)\cap M}\Big)+\underbrace{m}_{\in\operatorname{Supp}(h_{\Gamma})}>\operatorname{Min}_{\Delta}(n_{\Gamma})$
a contradiction. ∎
Lemma 8.6.
Assume that
$h_{\Gamma}\cdot g_{\Gamma}(f)\in L(\Delta)$
for some facet $\Gamma$. Then
$x^{m}\cdot g_{\Gamma}(f)\in L(\Delta)\quad\forall\,m\in\operatorname{Supp}(h_{\Gamma})$ (23)
Proof.
This is simple: Assume to the contrary that $x^{m}\cdot g_{\Gamma}(f)\notin L(\Delta)$, say that this polynomial jumps out of a facet $\Gamma_{1}$. Then choose $n\in N$ suitable and find a vertex $v\in Vert(\Gamma_{1})$ such that $m+v\notin\Delta\cap M$ and this vector remains left in $\operatorname{Supp}(h_{\Gamma}\cdot g_{\Gamma}(f))$. ∎
Define the „compensable sets“
$P_{\Gamma}:=$ $\{(m,v)\in\operatorname{Supp}(h_{\Gamma})\times\operatorname{Supp}(g_{\Gamma}(f))\mid m+v\notin\Delta\cap M\}$ (24)
[…]
For simplicity assume in the following $k=1$. Choose one aribitrary distinguished facet $\Gamma_{1}$ and restrict to those $(m,v)\in P_{\Gamma}$ with $\langle m+v,n_{1}\rangle<\operatorname{Min}_{\Delta}(n_{1})$ and without restriction $\langle m+v,n_{1}\rangle=\textrm{ minimal}$ for some fixed facet $\Gamma_{1}$ with inner facet normal $n_{1}$. We adopt the obvious notation $P_{i}$, $n_{i}$ and $h_{i}$.
Lemma 8.7.
In the above situation we may assume $\operatorname{Supp}(h_{1})=\emptyset$
Proof.
Proof: Else there is $(m,v)\in P_{1}$. Distinguish
  1. (a)
    $\exists m\in\operatorname{Supp}(h_{1})\cap\operatorname{Int}(\Gamma_{1})\cap M$ or
  2. (b)
    $\exists m\in\operatorname{Supp}(h_{1})\cap\operatorname{Int}(\Delta)\cap M$.
To (a): Choose $v\in Vert(\Delta)\setminus Vert(\Gamma_{1})$ with $\langle v,n_{1}\rangle$ minimal. Then $(m,v)\in P_{1}$,   $\langle v,n_{1}\rangle<0$ by (22) and there is $(m_{s},v_{s})\in P_{s}$ with
$m+v=m_{s}+v_{s}\Rightarrow\langle m_{s},n_{1}\rangle<0$ (25)
for some $s\geq 2$ (by Lemma 8.6) and
$\langle m,n_{1}\rangle=\operatorname{ord}_{\Delta}(n_{1})<\langle m_{s},n_{1}\rangle.$ (26)
Thus $v_{s}\in\Gamma_{1}$, else (25) gives a contradiction to the minimality of $\langle v,n_{1}\rangle$.
[…]
We differ
  1. (aenumi)
    $m_{s}\in\operatorname{Int}(\Delta)\cap M$ or
  2. (benumi)
    $m_{s}\in\operatorname{Int}(\Gamma_{s})\cap M$.
To (aenumi):
$\langle m+m_{s},n_{1}\rangle=\langle m+v,n_{1}\rangle<\operatorname{ord}_{\Delta}(n_{1})$
but
$\underbrace{m}_{\in\operatorname{Supp}(h_{1})}+\underbrace{m_{s}}_{\in\operatorname{Int}(\Delta)\cap M}\in R_{\Gamma_{1}}$
by (8.4), a contradiction.
To (benumi): Here
$\langle\operatorname{Int}(\Delta)\cap M,n_{s}\rangle\geq 0,\quad\quad\langle\operatorname{Int}(\Delta)\cap M,n_{1}\rangle\geq 0,$ (27)
since else
$\big(\operatorname{Int}(\Delta)\cap M\big)+m_{i}\notin\Delta\cap M,\quad\textrm{for }i=1,s.$
If the equation for $m_{s}$ is an equality, then apparently $\operatorname{Int}(\Delta)\cap M$ is contained in an hyperplane, a contradiction. Take $\tilde{m}\in\operatorname{Int}(\Delta)\cap M$ with $\langle\tilde{m},n_{s}\rangle>0$. If $\langle\tilde{m},n_{1}\rangle=0$ then
$\tilde{m}+m_{s}\in\operatorname{Int}(\Delta)\cap M$ (28)
and
$\langle m+\tilde{m}+m_{s},n_{1}\rangle=\langle m+m_{s},n_{1}\rangle<\operatorname{Min}_{\Delta}(n_{1})$
but $m+\tilde{m}+m_{s}\in R_{\Gamma_{1}}$, a contradiction. If $\langle\tilde{m},n_{1}\rangle>0$ then
$\langle\tilde{m}+k\cdot m,n_{1}\rangle=0\textrm{ for some }k\geq 1$
Replace $\tilde{m}$ by $\tilde{m}+k\cdot m$ and continue as above.
To (benumi): Take $k:=k(m)\in\mathbb{N}_{\geq 1}$. There is $(m_{s},v_{s})\in P_{s}$ with
$\underbrace{m}_{\in\operatorname{Int}(\Delta)\cap M}+\underbrace{v}_{\langle v,n_{1}\rangle=min.,v\notin\Gamma_{1}}=\underbrace{m_{s}}_{\langle m_{s},n_{1}\rangle<0\textrm{ since }(m_{s},v_{s})\in P_{1}}+v_{s}.$
If $v_{s}\in\Gamma_{1}$, choose $\tilde{v}\in\Gamma_{1}$ such that $(m_{s},\tilde{v})\in P_{1}$
$\langle m_{s}+\tilde{v},n_{1}\rangle<\langle m_{s}+v_{s},n_{1}\rangle=\textrm{min.}$
contradiction. Thus $v_{s}\in\Gamma_{1}$. Then
$\underbrace{\langle m,n_{1}\rangle}_{\textrm{minimal},<0}+\underbrace{\langle v,n_{1}\rangle}_{>\textrm{ min}}$ $=\langle m+v,n_{1}\rangle$
$=\langle m_{s}+v_{s},n_{1}\rangle$
$=\underbrace{\langle m_{s},n_{1}\rangle}_{<0}+\underbrace{\langle v_{s},n_{1}\rangle}_{=\textrm{ min.}}$
$\Rightarrow 0>\langle m_{s},n_{1}\rangle>\langle m,n_{1}\rangle$. A contradiction to (22). ∎
By the above Lemma we may replace $f\mapsto f_{|{\Gamma_{1}}}$ and thus assume $\operatorname{Supp}(f)\subset\Gamma_{1}\cap M$. We get
$\sum\limits_{\Gamma}h_{\Gamma}\cdot g_{\Gamma}(f_{|{\Gamma_{1}}})=0$ (29)
and by the assumptions $\operatorname{Supp}(h_{\Gamma})$ is contained in a plane $H_{n_{1},-l}$ for some $l$. Say $n_{1}=e_{n}=(0,...,1)$. Consider the Laurent polynomials $F_{0},...,F_{n+1}$: The last $(F_{n+1})_{|{\Gamma_{1}}}$ is proportional to $x_{0}\cdot f_{|{\Gamma_{1}}}$, that is to $F_{0}$. Further the operations $f\mapsto f_{|{\Gamma_{1}}}$ and $f\mapsto F_{i}$ commute. From the defining equations of the $g_{\Gamma}(f)$ in Construction (3.2) and the equation (29) we get an algebraic depependence between $F_{0},...,F_{n}$ and, $f_{|{\Gamma_{1}}}$ remaining nondegenerate, it results $h_{\Gamma}\equiv 0$ by Theorem (2.4), as desired.
Corollary 8.8.
Under the assumption of theorem (8.4) we have
$\ker(d\phi_{f})=\ker(d\phi_{f}^{1}).$
Proof.
We show
$\ker(d\phi_{f}^{k})\subset\ker(d\phi_{f}^{k-1})\quad k=2,...,n$
$\ker(d\phi_{f}^{n-1})=\ker(d\phi_{f}^{1}).$
To the first point: Inductively it is enough to show
$\ker(d\phi_{f}^{n})\subset\ker(d\phi_{f}^{n-1}).$
We show by contradiction given $g_{\Gamma}(f)\cdot x^{w}\notin\ker(d\phi_{f}^{n-1})$ then $g_{\Gamma}(f)\cdot x^{w}\notin\ker(\phi_{f}^{n})$: There is $v\in\operatorname{Int}((n-1)\cdot\Delta)$ with one of the following
[…]
$\langle w+(v+\tilde{v}),n_{\tilde{\Gamma}}\rangle$ $\leq n\cdot\operatorname{Min}_{\Delta}(n_{\tilde{\Gamma}})+\langle\tilde{v},n_{\tilde{\Gamma}}\rangle$
[…]
Thus $w\notin\ker(d\phi_{f}^{n})$, a contradiction.
To the second point: Choose $\tilde{v}\in Vert(\Gamma)$, then $v+\tilde{v}\in\operatorname{Int}(n\cdot\Delta)\cap M$ for all $v$ and
$\langle w+(v+\tilde{v}),n_{\Gamma}\rangle$ $<n\cdot\operatorname{Min}_{\Delta}(n_{\Gamma})+\langle\tilde{v},n_{\Gamma}\rangle$
$=(n+1)\cdot\operatorname{Min}_{\Delta}(n_{\Gamma})$
Thus $w\notin\ker(d\phi_{f}^{n})$, contradiction.
Last but not least we have
$\operatorname{Hom}(R_{Int,f}^{n-1},R_{Int,f}^{n})$ $\cong(R_{Int,f}^{n-1})^{*}\otimes R_{Int,f}^{n}$
$R_{Int,f}^{2}\otimes(R_{Int,f}^{1})^{*}$
$\operatorname{Hom}(R_{Int,f}^{1},R_{Int,f}^{2})$
and $d\phi_{f}^{1}$ corresponds to $d\phi_{f}^{n-1}$ under these isomorphisms. ∎
Example 8.9.
Take $\mathbb{P}=\mathbb{P}^{n}$, where $n\geq 3$ and $Y_{f}$ a nondegenerate (not exactly the same as smooth) hypersurface in $\mathbb{P}^{n}$ of degree $\geq n+2$. Then the assumptions of theorem (8.4) are met and the Kodaira-Spencer map is surjective (see [Voi03, Lemma 6.15]). Thus we get a proof of the ITT for such projective hypersurfaces.
Corollary 8.10.
Given $n=3$ assume $\operatorname{Int}(\Delta)\cap M\nsubseteq\textrm{plane}$. Then
$\ker(\Phi_{f|{Im\,\kappa_{f}}})=\{0\},$
that is $Y_{f}$ fulfills the ITT for $\mathcal{X}\rightarrow B$.
Proof.
Let $f\in B$ and $g\in\ker(\phi_{f}^{1})$. By theorem 8.4 we may assume that
$g=g_{\Gamma}(f)\cdot x^{w},$
where
$w\in R_{\Gamma}\cap M$
for some $\Gamma\leq\Delta$ and $\langle w,n_{\Gamma}\rangle=-k\leq 0$, $\langle w,n_{\Gamma^{\prime}}\rangle\geq 0$ for $\Gamma^{\prime}\neq\Gamma$. There are two simple cases
  • $k=0$: Then $w=(0,0,0)$ and $g=g_{\Gamma}(f)\equiv 0\in R_{f}^{1}$.
  • k=1k=-1: Then $w\in-R(N,\Sigma_{\Delta})$ and $g\in\ker(\kappa_{f})$.
Else $k\leq-2$ and there is no lattice point $w^{\prime}\in\operatorname{Int}(\Delta)\cap M$ with
$0>\langle w^{\prime},n_{\Gamma}\rangle>\langle w,n_{\Gamma}\rangle.$
Choose $x:=(0,0,0)$, and $y,z\in\operatorname{Int}(\Delta)\cap M$ not lying on a line and with $|\langle x,y,z\rangle\cap M|=3$. Up to adding some multiple of $w$ to $x,y,z,w$ we may assume that
$x,y,z\in H_{n_{\Gamma},0},\quad w\in H_{n_{\Gamma},-l_{\Gamma}}$
and
$|\langle x,y,z,w\rangle\cap M|=4.$
Here we implicitly used the formula of theorem (8.4). By White’s Theorem ([Oda88, p.34, the Terminal Lemma]) we may assume
$\{x\}=\begin{pmatrix}0\\ 0\\ 0\end{pmatrix},\{y,z,w\}=\{\begin{pmatrix}1\\ 0\\ 0\end{pmatrix},\,\begin{pmatrix}0\\ 0\\ 1\end{pmatrix},\,\begin{pmatrix}1\\ p\\ q\end{pmatrix}\}$ (30)
where $0<p<q$ and $\gcd(p,q)=1$. Besides apparently for $Q:=\langle x,y,z,w,y+w,z+w\rangle$ we have $|Q\cap M|=6$ since we translate the triangle $\langle x,y,z\rangle$ by $w$ (compare Lemma 8.5). The point is that the assumptions and this explicit representation force $|Q\cap M|\geq 7$, a contradiction:
If $w=(1,0,0)$ then it follows easily from $\langle w,n_{\Gamma}\rangle\leq-2$ that $p\geq 2$.
$Q=\langle\begin{pmatrix}0\\ 0\\ 0\end{pmatrix},\begin{pmatrix}1\\ 0\\ 0\end{pmatrix},\begin{pmatrix}0\\ 0\\ 1\end{pmatrix},\begin{pmatrix}1\\ p\\ q\end{pmatrix},\begin{pmatrix}1\\ 0\\ 1\end{pmatrix},\begin{pmatrix}2\\ p\\ q\end{pmatrix}\rangle$.
Distinguish if $q$ is even and $p$ is odd let $q=r\cdot p+s$ with $r,s\in\mathbb{N}_{\geq 1}$ and $0<s<p$. Then
$\frac{1}{p}\cdot\Big(\begin{pmatrix}1\\ p\\ q\end{pmatrix}+(p-s)\cdot\begin{pmatrix}1\\ 0\\ 1\end{pmatrix}+(s-1)\cdot\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}\Big)=\begin{pmatrix}1\\ 1\\ r+1\end{pmatrix}\in Q\cap M$
If $q$ is odd and $p$ is even then
$\frac{1}{2}\cdot\Big(\begin{pmatrix}2\\ p\\ q\end{pmatrix}+(q-p)\cdot\begin{pmatrix}0\\ 0\\ 1\end{pmatrix}=\begin{pmatrix}1\\ \frac{p}{2}\\ \frac{q+1}{2}\end{pmatrix}\in Q\cap M$
If $w=(0,0,1)$ then again $p\geq 2$ and
$Q=\langle\begin{pmatrix}0\\ 0\\ 0\end{pmatrix},\begin{pmatrix}1\\ 0\\ 0\end{pmatrix},\begin{pmatrix}0\\ 0\\ 1\end{pmatrix},\begin{pmatrix}1\\ p\\ q\end{pmatrix},\begin{pmatrix}1\\ 0\\ 1\end{pmatrix},\begin{pmatrix}1\\ p\\ q+1\end{pmatrix}\rangle$.
If $q$ is even and $p$ is odd again $(1,1,r+1)\in Q\cap M$. If $q$ is odd and $p$ is even then $(1,\frac{p}{2},\frac{q+1}{2})\in Q\cap M$ as above.
Finally if $w=(1,p,q)$ then $p\geq 2$ and
$Q=\langle\begin{pmatrix}0\\ 0\\ 0\end{pmatrix},\begin{pmatrix}1\\ 0\\ 0\end{pmatrix},\begin{pmatrix}0\\ 0\\ 1\end{pmatrix},\begin{pmatrix}1\\ p\\ q\end{pmatrix},\begin{pmatrix}2\\ p\\ q\end{pmatrix},\begin{pmatrix}1\\ p\\ q+1\end{pmatrix}\rangle$.
If $q$ is even and $p$ is odd let $r:=q\,\mod\,p$. There are $\lambda,\mu\geq 0$ with
$\lambda+\mu=r,\quad\lambda\cdot q+\mu\cdot(q+1)=l\cdot p,\quad l\in\mathbb{Z}.$
Then
$\frac{1}{p}\cdot\big((p-r)\cdot\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}+\lambda\cdot\begin{pmatrix}1\\ p\\ q\end{pmatrix}+\mu\cdot\begin{pmatrix}1\\ p\\ q+1\end{pmatrix}\big)=\begin{pmatrix}1\\ r\\ l\end{pmatrix}\in Q\cap M.$
Remark 8.11.
Note that we consider some known counterexamples to the ITT in ([Gie23]), namely special Kanev and Todorov surfaces in toric $3$-folds, where $\operatorname{Int}(\Delta)\cap M=\{(0,0,0)\}$, but $\dim\,F(\Delta)=3$ . Whereas the dimension of the kernel of $\kappa_{f}$ turn out to be independent of $f\in B$ the dimension of thekernel of $d\mathcal{P}_{B,f}$ in general depends on $f\in B$. This is because we force $g_{\Gamma}(f)\cdot x^{w}\in L(\Delta)$ which might very well depend on $f\in B$.